OpenStudy (anonymous):

for the following function g(x)=2x^2-4 what is g(-2)

OpenStudy (anonymous):

-20

OpenStudy (anonymous):

2(-2)^2-4=20 so g(x)=20

OpenStudy (anonymous):

20 is not one of my choices

jimthompson5910 (jim_thompson5910):

g(-2)=2(-2)^2-4=2(4)-4=8-4=4 So g(-2) = 4

OpenStudy (anonymous):

i am thinking -12

OpenStudy (anonymous):

jim is right, i did the equation in the wrong order of operations, lol :(